# Optimal frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v1

@@ 3.1. The Relationship between Tight Frames and Convex Polygons in $\mathbb{R}^{2}$
Let $\mathcal{A}=\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\subset\mathbb{R}^{d}$. If there exist two distinct integers $i,j\in[m]$ such that $\bm{a}_{i}$ is a scalar multiple of $\bm{a}_{j}$, i.e., $\bm{a}_{i}=\lambda\bm{a}_{j}$ for some $\lambda\in\mathbb{R}$, then $\mathcal{A}$ is called reducible; otherwise, $\mathcal{A}$ is called irreducible.
@@
Assume that $\mathcal{A}=\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}\subset\mathcal{T}$, where $\mathcal{T}$ is defined in (1.9). Let $E=\{\bm{e}_{1},\dotsc,\bm{e}_{m}\}$, where each $\bm{e}_{j}=f(\bm{a}_{j})$, where $f$ is defined in (1.10). Then we have the following results:
@@
    The set $\mathcal{A}$ is a tight frame if and only if $\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}$.
@@
    Assume that the set $\mathcal{A}$ forms a tight frame. Let $\bm{A}=[\bm{a}_{1},\dotsc,\bm{a}_{m}]^{T}$. If $\mathcal{A}$ is irreducible, then
    $\beta_{\bm{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}},$
    where $P$ is the strictly convex $m$-gon defined in (iii). If $\mathcal{A}$ is reducible, then the condition number of $\bm{A}$ is not the minimum value attainable among all $m\times 2$ matrices.
@@
For clarity, we write $\bm{a}_{j}=t_{j}(\cos{\phi_{j}},\sin{\phi_{j}})^{T}$ for each $j\in[m]$, where $t_{j}\geq 0$ and $\phi_{j}\in(0,\pi]$.
(i) We prove the equivalence in two directions. Assume that the set $\mathcal{A}$ is irreducible. First, we show that vectors in $E$ are nonzero. It is important to note that $\bm{e}_{j}=f(\bm{a}_{j})$ is the zero vector if and only if $\bm{a}_{j}$ is the zero vector. Since $\mathcal{A}$ is irreducible, by definition, it does not contain the zero vector. Consequently, given this property of $f$, the set $E$ also does not contain the zero vector. Next, we show that vectors in $E$ have pairwise distinct directions. Suppose, for contradiction, that there exist distinct vectors $\bm{e}_{i}=f(\bm{a}_{i})$ and $\bm{e}_{j}=f(\bm{a}_{j})$ (with $i\neq j$) that have the same direction. This implies that their associated directional parameters are equal, i.e., $\phi_{j}=\phi_{i}$. This condition is, in turn, equivalent to $\bm{a}_{j}=\lambda\bm{a}_{i}$ for some nonzero scalar $\lambda\in\mathbb{R}$. However, this directly contradicts the definition of an irreducible set $\mathcal{A}$, which requires that no two distinct vectors are scalar multiples of each other. Hence, we conclude that if $\mathcal{A}$ is irreducible, then the vectors in $E$ are nonzero and have pairwise distinct directions.
@@
(ii) We define $a=\sum_{j=1}^{m}t_{j}^{2}\cos^{2}{\phi_{j}}$, $b=\sum_{j=1}^{m}t_{j}^{2}\sin^{2}{\phi_{j}}$ and $c=\sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}$. A direct computation shows that $\sum_{j=1}^{m}\bm{e}_{j}=(a-b,2c)^{T}$ and
$\sum_{j=1}^{m}\bm{a}_{j}\bm{a}_{j}^{T}=\begin{pmatrix}\sum_{j=1}^{m}t_{j}^{2}\cos^{2}{\phi_{j}}&\sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}\\ \sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}&\sum_{j=1}^{m}t_{j}^{2}\sin^{2}{\phi_{j}}\\ \end{pmatrix}=\begin{pmatrix}a&c\\ c&b\\ \end{pmatrix}$. (3.1)
Therefore, $\{\bm{a}_{1},\dots,\bm{a}_{m}\}$ is a tight frame if and only if $a=b$ and $c=0$, which is equivalent to $\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}$.
(iii) Assume that $\{\bm{a}_{1},\dots,\bm{a}_{m}\}$ is an irreducible tight frame. It follows from (i) and (ii) that $\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}$ and $\bm{e}_{1},\dots,\bm{e}_{m}$ have pairwise distinct directions. Since $\bm{e}_{j}\neq\bm{0}$ for all $j\in[m]$, according to Corollary 2.1, there exists a strictly convex $m$-gon P{P} with the edge set $E$.
Assume that $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$. By Corollary 2.1 we have $\sum_{j=1}^{m}\bm{e}_{j}=\bm{0}$, and $\bm{e}_{1},\dots,\bm{e}_{m}$ are nonzero vectors with pairwise distinct directions. If follows from (i) and (ii) that $\{\bm{a}_{1},\dots,\bm{a}_{m}\}$ is an irreducible tight frame.
@@
$\beta_{\bm{A}}=\left(1-\frac{1}{C}\max\limits_{\bm{x},\bm{y}\in\mathbb{S}^{1},\langle\bm{x},\bm{y}\rangle=0}\sum_{j=1}^{m}\left|\bm{x}^{T}\bm{a}_{j}\bm{a}_{j}^{T}\bm{y}\right|\right)^{-\frac{1}{2}},$ (3.2)
where $C=\frac{1}{2}\sum_{j=1}^{m}{\|\bm{a}_{j}\|_{2}^{2}}=\frac{1}{2}\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}$. Let $\bm{x}=(\cos\theta,\sin\theta)^{T}$, $\bm{y}=(-\sin\theta,\cos\theta)^{T}$ and $\bm{u}=(\sin 2\theta,\cos 2\theta)^{T}$, where $\theta\in[0,2\pi)$. A direct calculation shows that for each $j\in[m]$
$\left|\bm{x}^{T}\bm{a}_{j}\bm{a}_{j}^{T}\bm{y}\right|=\frac{t_{j}^{2}}{2}|\sin(2\theta-2\phi_{j})|=\frac{1}{2}|\langle\bm{e}_{j},\bm{u}\rangle|.$
@@
$\max\limits_{\bm{x},\bm{y}\in\mathbb{S}^{1},\langle\bm{x},\bm{y}\rangle=0}\sum_{j=1}^{m}\left|\bm{x}^{T}\bm{a}_{j}\bm{a}_{j}^{T}\bm{y}\right|=\max\limits_{\theta\in[0,2\pi)}\sum_{j=1}^{m}\frac{t_{j}^{2}}{2}|\sin(2\theta-2\phi_{j})|=\frac{1}{2}\cdot\max\limits_{\bm{u}\in\mathbb{S}^{1}}\sum_{j=1}^{m}|\langle\bm{e}_{j},\bm{u}\rangle|.$ (3.3)
Substituting (3.3) and $C=\frac{1}{2}\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}$ into (3.2), we obtain
$\beta_{\bm{A}}=\left(1-\frac{\max\limits_{\bm{u}\in\mathbb{S}^{1}}\sum_{j=1}^{m}|\langle\bm{e}_{j},\bm{u}\rangle|}{\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}}\right)^{-\frac{1}{2}}.$ (3.4)
If $\mathcal{A}$ is irreducible, it follows from (iii) that $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$. Combining (3.4) with Lemma 2.4, we obtain $\beta_{\bm{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}}$.
We next prove that the condition number of $\bm{A}$ is not minimal if $\mathcal{A}$ is reducible. Assume $\mathcal{A}$ is reducible. Then the set $E$ contains zero vectors or vectors that have the same direction. By deleting zero vectors from $E$ and replacing all vectors in the same direction with their sum, we obtain a new set $\widehat{E}=\{\widehat{\bm{e}}_{1},\dotsc,\widehat{\bm{e}}_{k}\}$ with $k<m$. The vectors in $\widehat{E}$ are nonzero and have pairwise distinct directions. Note that $\sum_{j=1}^{k}\widehat{\bm{e}}_{j}=\sum_{j=1}^{m}\bm{e}_{j}=0$. By Corollary 2.1, there exists a strictly convex $k$-gon $\widehat{P}\in\mathcal{P}_{k}$ with the edge set $\widehat{E}$. Note that $\sum_{j=1}^{m}\|\bm{e}_{j}\|_{2}=\sum_{j=1}^{k}\|\widehat{\bm{e}}_{j}\|_{2}$ and $\sum_{j=1}^{m}|\langle\bm{e}_{j},\bm{u}\rangle|=\sum_{j=1}^{k}|\langle\widehat{\bm{e}}_{j},\bm{u}\rangle|$ for any $\bm{u}\in\mathbb{S}^{1}$. Combining with (3.4) and Lemma 2.4, we have
$\beta_{\bm{A}}=\frac{1}{\sqrt{1-2\cdot r(\widehat{P})}}$. (3.5)
Let $Q\in\mathcal{P}_{m}$ be an optimal strictly convex $m$-gon. Denote the edge set of $Q$ by $\{\bm{q}_{1},\ldots,\bm{q}_{m}\}$. Let $\bm{B}=[f^{-1}(\bm{q}_{1}),\ldots,f^{-1}(\bm{q}_{m})]^{T}\in\mathbb{R}^{m\times 2}$, where $f(\cdot)$ is defined in (1.10). According to (iii), the set $\{f^{-1}(\bm{q}_{1}),\ldots,f^{-1}(\bm{q}_{m})\}$ forms an irreducible tight frame. Then, we have
$\beta_{\bm{B}}=\frac{1}{\sqrt{1-2\cdot r({Q})}}.$ (3.6)
@@
Combining with (3.5) and (3.6), we have $\beta_{\bm{A}}>\beta_{\bm{B}}$. Therefore, the condition number of $\bm{A}$ is not minimal. This completes the proof.
@@
(i) Assume that $\mathcal{A}$ is an optimal tight frame, i.e., the matrix $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{m}]^{T}$ has the minimal condition number and its rows form a tight frame. It follows from Lemma 3.1 (iv) that $\mathcal{A}$ is irreducible. Then Lemma 3.1 (iii) shows that the set $\{f(\bm{a}_{1}),\dotsc,f(\bm{a}_{m})\}$ is the edge set of a strictly convex $m$-gon ${P}\subset\mathbb{R}^{2}$, and we have $\beta_{{\bm{A}}}=\frac{1}{\sqrt{1-2\cdot r(P)}}$. We prove that $P$ is optimal by contradiction. Suppose, to the contrary, that $P$ is not optimal. Then there exists a strictly convex $m$-gon ${Q}\subset\mathbb{R}^{2}$ such that $r(Q)<r(P)$. Denote the edge vectors set of $Q$ by $\{\bm{q}_{1},\ldots,\bm{q}_{m}\}$. Let $\bm{B}=[f^{-1}(\bm{q}_{1}),\ldots,f^{-1}(\bm{q}_{m})]^{T}\in\mathbb{R}^{m\times 2}$. By Lemma 3.1 (iii) we have $\beta_{\bm{B}}=\frac{1}{\sqrt{1-2\cdot r({Q})}}$. Since $r(Q)<r(P)$, we have $\beta_{\bm{B}}<\beta_{\bm{A}}$. This contradicts with the fact that $\bm{A}$ has the minimal condition number. Hence, $P$ is optimal.
(ii) Assume that $\{f(\bm{a}_{1}),\dotsc,f(\bm{a}_{m})\}$ is the edge set of an optimal $m$-gon $P\in\mathcal{P}_{m}$. Then $P$ is strictly convex. According to Lemma 3.1 (iii) and (iv), the set $\{\bm{a}_{1},\dotsc,\bm{a}_{m}\}$ is an irreducible tight frame, and we have $\beta_{{\bm{A}}}=\frac{1}{\sqrt{1-2\cdot r(P)}}$. We prove that $\bm{A}$ has the minimal condition number by contradiction. Assume that $\beta_{\bm{B}}<\beta_{\bm{A}}$, where $\bm{B}\in\mathbb{R}^{m\times 2}$ is a matrix with the minimal condition number and its rows form a tight frame. By Lemma 3.1 (iii), the rows of $\bm{B}$ is irreducible, and there is a strictly convex $m$-gon ${Q}\subset\mathbb{R}^{2}$ such that $\beta_{{\bm{B}}}=\frac{1}{\sqrt{1-2\cdot r(Q)}}$. Since $\beta_{\bm{B}}<\beta_{\bm{A}}$, we have $r(Q)<r(P)$. This contradicts with $P$ being an optimal convex $m$-gon. Hence, $\bm{A}$ has the minimal condition number, and its rows form a tight frame.
@@
$\min_{\bm{A}\in\mathbb{R}^{m\times 2}}\beta_{\bm{A}}=\min_{\bm{A}\in\mathbb{R}^{m\times 2},\bm{A}^{T}\bm{A}=\bm{I}_{2}}\beta_{\bm{A}}.$
Let $\bm{B}\in\mathbb{R}^{m\times 2}$ be such that $\bm{B}^{T}\bm{B}=\bm{I}_{2}$ and the condition number of $\bm{B}$ is minimal. By Theorem 1.2, there is an optimal convex $m$-gon ${P}\subset\mathbb{R}^{2}$ such that
$\min_{\bm{A}\in\mathbb{R}^{m\times 2}}\beta_{\bm{A}}=\beta_{\bm{B}}=\frac{1}{\sqrt{1-2\cdot r(P)}}.$ (3.7)
@@
$\min_{\bm{A}\in\mathbb{R}^{m\times 2}}\beta_{\bm{A}}=\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$
@@
$\min_{\bm{A}\in\mathbb{R}^{m\times 2}}\beta_{\bm{A}}>\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$
(iii) Recall that $P_{m}$ is the regular $m$-gon. Theorem 1.1 (iii) shows that $r(P_{m})=r(P)$ if and only if $m$ is odd, and $P_{m}$ is uniquely optimal if and only if $m$ is a prime. According to Theorem 1.2, the matrix constructed from $P_{m}$ corresponds to the harmonic frame $\bm{E}_{m}$; in particular, $\bm{E}_{m}$ consists of the vectors $f^{-1}(\bm{e}_{j})$ obtained from the edge vectors $\bm{e}_{j}$ of $P_{m},j=1,\ldots,m$. Combining Theorem 1.2 and Theorem 1.1, we conclude that $\bm{E}_{m}$ has the minimal condition number if and only if $m$ is odd. Moreover, $\bm{E}_{m}$ is the unique matrix with minimal condition number if and only if $m$ is a prime. ∎

### v2

@@ 3.1. The relationship between tight frames and convex polygons in $\mathbb{R}^{2}$
Let $\mathcal{A}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{R}^{d}$. If there exist two distinct integers $i,j\in[m]$ such that $\boldsymbol{a}_{i}$ is a scalar multiple of $\boldsymbol{a}_{j}$, i.e., $\boldsymbol{a}_{i}=\lambda\boldsymbol{a}_{j}$ for some $\lambda\in\mathbb{R}$, then $\mathcal{A}$ is called reducible; otherwise, $\mathcal{A}$ is called irreducible.
@@
Assume that $\mathcal{A}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathcal{T}$, where $\mathcal{T}$ is defined in (1.9). Let $E=\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}$, where each $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})$, where $f$ is defined in (1.10). Then we have the following results:
@@
    The set $\mathcal{A}$ is a tight frame if and only if $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$.
@@
    Assume that the set $\mathcal{A}$ forms a tight frame. Let $\boldsymbol{A}=[\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}]^{T}$. If $\mathcal{A}$ is irreducible, then
    $\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}},$
    where $P$ is the strictly convex $m$-gon defined in (iii). If $\mathcal{A}$ is reducible, then the condition number of $\boldsymbol{A}$ is not the minimum value attainable among all $m\times 2$ matrices.
@@
For clarity, we write $\boldsymbol{a}_{j}=t_{j}(\cos{\phi_{j}},\sin{\phi_{j}})^{T}$ for each $j\in[m]$, where $t_{j}\geq 0$ and $\phi_{j}\in(0,\pi]$.
(i) We prove the equivalence in two directions. Assume that the set $\mathcal{A}$ is irreducible. First, we show that vectors in $E$ are nonzero. It is important to note that $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})$ is the zero vector if and only if $\boldsymbol{a}_{j}$ is the zero vector. Since $\mathcal{A}$ is irreducible, by definition, it does not contain the zero vector. Consequently, given this property of $f$, the set $E$ also does not contain the zero vector. Next, we show that vectors in $E$ have pairwise distinct directions. Suppose, for contradiction, that there exist distinct vectors $\boldsymbol{e}_{i}=f(\boldsymbol{a}_{i})$ and $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})$ (with $i\neq j$) that have the same direction. This implies that their associated directional parameters are equal, i.e., $\phi_{j}=\phi_{i}$. This condition is, in turn, equivalent to $\boldsymbol{a}_{j}=\lambda\boldsymbol{a}_{i}$ for some nonzero scalar $\lambda\in\mathbb{R}$. However, this directly contradicts the definition of an irreducible set $\mathcal{A}$, which requires that no two distinct vectors are scalar multiples of each other. Hence, we conclude that if $\mathcal{A}$ is irreducible, then the vectors in $E$ are nonzero and have pairwise distinct directions.
@@
(ii) We define
$a=\sum_{j=1}^{m}t_{j}^{2}\cos^{2}{\phi_{j}},\quad b=\sum_{j=1}^{m}t_{j}^{2}\sin^{2}{\phi_{j}},\quad\text{and}\quad c=\sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}.$
A direct computation shows that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=(a-b,2c)^{T}$ and
$\sum_{j=1}^{m}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{T}=\begin{pmatrix}\sum_{j=1}^{m}t_{j}^{2}\cos^{2}{\phi_{j}}&\sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}\\ \sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}&\sum_{j=1}^{m}t_{j}^{2}\sin^{2}{\phi_{j}}\\ \end{pmatrix}=\begin{pmatrix}a&c\\ c&b\\ \end{pmatrix}$. (3.1)
Therefore, $\{\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{m}\}$ is a tight frame if and only if $a=b$ and $c=0$, which is equivalent to $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$.
(iii) Assume that $\{\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{m}\}$ is an irreducible tight frame. It follows from (i) and (ii) that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$ and $\boldsymbol{e}_{1},\dots,\boldsymbol{e}_{m}$ have pairwise distinct directions. Since $\boldsymbol{e}_{j}\neq\boldsymbol{0}$ for all $j\in[m]$, according to Corollary 2.1, there exists a strictly convex $m$-gon P{P} with the edge set $E$.
Assume that $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$. By Corollary 2.1 we have $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$, and $\boldsymbol{e}_{1},\dots,\boldsymbol{e}_{m}$ are nonzero vectors with pairwise distinct directions. If follows from (i) and (ii) that $\{\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{m}\}$ is an irreducible tight frame.
@@
$\beta_{\boldsymbol{A}}=\left(1-\frac{1}{C}\max\limits_{\boldsymbol{x},\boldsymbol{y}\in\mathbb{S}^{1},\langle\boldsymbol{x},\boldsymbol{y}\rangle=0}\sum_{j=1}^{m}\left|\boldsymbol{x}^{T}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{T}\boldsymbol{y}\right|\right)^{-\frac{1}{2}},$ (3.2)
where $C=\frac{1}{2}\sum_{j=1}^{m}{\|\boldsymbol{a}_{j}\|_{2}^{2}}=\frac{1}{2}\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}$. Let
$\boldsymbol{x}=(\cos\theta,\sin\theta)^{T},\quad\boldsymbol{y}=(-\sin\theta,\cos\theta)^{T},\quad\text{and}\quad\boldsymbol{u}=(\sin 2\theta,\cos 2\theta)^{T},$
where $\theta\in[0,2\pi)$. A direct calculation shows that for each $j\in[m]$
$\left|\boldsymbol{x}^{T}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{T}\boldsymbol{y}\right|=\frac{t_{j}^{2}}{2}|\sin(2\theta-2\phi_{j})|=\frac{1}{2}|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|.$
@@
$\max\limits_{\boldsymbol{x},\boldsymbol{y}\in\mathbb{S}^{1},\langle\boldsymbol{x},\boldsymbol{y}\rangle=0}\sum_{j=1}^{m}\left|\boldsymbol{x}^{T}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{T}\boldsymbol{y}\right|=\max\limits_{\theta\in[0,2\pi)}\sum_{j=1}^{m}\frac{t_{j}^{2}}{2}|\sin(2\theta-2\phi_{j})|=\frac{1}{2}\cdot\max\limits_{\boldsymbol{u}\in\mathbb{S}^{1}}\sum_{j=1}^{m}|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|.$ (3.3)
Substituting (3.3) and $C=\frac{1}{2}\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}$ into (3.2), we obtain
$\beta_{\boldsymbol{A}}=\left(1-\frac{\max\limits_{\boldsymbol{u}\in\mathbb{S}^{1}}\sum_{j=1}^{m}|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|}{\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}}\right)^{-\frac{1}{2}}.$ (3.4)
If $\mathcal{A}$ is irreducible, it follows from (iii) that $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$. Combining (3.4) with Lemma 2.4, we obtain $\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}}$.
We next prove that the condition number of $\boldsymbol{A}$ is not minimal if $\mathcal{A}$ is reducible. Assume $\mathcal{A}$ is reducible. Then the set $E$ contains zero vectors or vectors that have the same direction. By deleting zero vectors from $E$ and replacing all vectors in the same direction with their sum, we obtain a new set $\widehat{E}=\{\widehat{\boldsymbol{e}}_{1},\dotsc,\widehat{\boldsymbol{e}}_{k}\}$ with $k<m$. The vectors in $\widehat{E}$ are nonzero and have pairwise distinct directions. Note that $\sum_{j=1}^{k}\widehat{\boldsymbol{e}}_{j}=\sum_{j=1}^{m}\boldsymbol{e}_{j}=0$. By Corollary 2.1, there exists a strictly convex $k$-gon $\widehat{P}\in\mathcal{P}_{k}$ with the edge set $\widehat{E}$. Note that $\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}=\sum_{j=1}^{k}\|\widehat{\boldsymbol{e}}_{j}\|_{2}$ and $\sum_{j=1}^{m}|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|=\sum_{j=1}^{k}|\langle\widehat{\boldsymbol{e}}_{j},\boldsymbol{u}\rangle|$ for any $\boldsymbol{u}\in\mathbb{S}^{1}$. Combining with (3.4) and Lemma 2.4, we have
$\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-2\cdot r(\widehat{P})}}$. (3.5)
Let $Q\in\mathcal{P}_{m}$ be an optimal strictly convex $m$-gon. Denote the edge set of $Q$ by $\{\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{m}\}$. Let $\boldsymbol{B}=[f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})]^{T}\in\mathbb{R}^{m\times 2}$, where $f(\cdot)$ is defined in (1.10). According to (iii), the set $\{f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})\}$ forms an irreducible tight frame. Then, we have
$\beta_{\boldsymbol{B}}=\frac{1}{\sqrt{1-2\cdot r({Q})}}.$ (3.6)
@@
Combining with (3.5) and (3.6), we have $\beta_{\boldsymbol{A}}>\beta_{\boldsymbol{B}}$. Therefore, the condition number of $\boldsymbol{A}$ is not minimal. This completes the proof.
@@
(i) Assume that $\mathcal{A}$ is an optimal tight frame, i.e., the matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}$ has the minimal condition number and its rows form a tight frame. It follows from Lemma 3.1 (iv) that $\mathcal{A}$ is irreducible. Then Lemma 3.1 (iii) shows that the set $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge set of a strictly convex $m$-gon ${P}\subset\mathbb{R}^{2}$, and we have $\beta_{{\boldsymbol{A}}}=\frac{1}{\sqrt{1-2\cdot r(P)}}$. We prove that $P$ is optimal by contradiction. Suppose, to the contrary, that $P$ is not optimal. Then there exists a strictly convex $m$-gon ${Q}\subset\mathbb{R}^{2}$ such that $r(Q)<r(P)$. Denote the edge vectors set of $Q$ by $\{\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{m}\}$. Let $\boldsymbol{B}=[f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})]^{T}\in\mathbb{R}^{m\times 2}$. By Lemma 3.1 (iii) we have $\beta_{\boldsymbol{B}}=\frac{1}{\sqrt{1-2\cdot r({Q})}}$. Since $r(Q)<r(P)$, we have $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$. This contradicts with the fact that $\boldsymbol{A}$ has the minimal condition number. Hence, $P$ is optimal.
(ii) Assume that $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge set of an optimal $m$-gon $P\in\mathcal{P}_{m}$. Then $P$ is strictly convex. According to Lemma 3.1 (iii) and (iv), the set $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}$ is an irreducible tight frame, and we have $\beta_{{\boldsymbol{A}}}=\frac{1}{\sqrt{1-2\cdot r(P)}}$. We prove that $\boldsymbol{A}$ has the minimal condition number by contradiction. Assume that $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$, where $\boldsymbol{B}\in\mathbb{R}^{m\times 2}$ is a matrix with the minimal condition number and its rows form a tight frame. By Lemma 3.1 (iii), the rows of $\boldsymbol{B}$ is irreducible, and there is a strictly convex $m$-gon ${Q}\subset\mathbb{R}^{2}$ such that $\beta_{{\boldsymbol{B}}}=\frac{1}{\sqrt{1-2\cdot r(Q)}}$. Since $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$, we have $r(Q)<r(P)$. This contradicts with $P$ being an optimal convex $m$-gon. Hence, $\boldsymbol{A}$ has the minimal condition number, and its rows form a tight frame.
@@
$\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}=\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2},\boldsymbol{A}^{T}\boldsymbol{A}=\boldsymbol{I}_{2}}\beta_{\boldsymbol{A}}.$
Let $\boldsymbol{B}\in\mathbb{R}^{m\times 2}$ be such that $\boldsymbol{B}^{T}\boldsymbol{B}=\boldsymbol{I}_{2}$ and the condition number of $\boldsymbol{B}$ is minimal. By Theorem 1.2, there is an optimal convex $m$-gon ${P}\subset\mathbb{R}^{2}$ such that
$\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}=\beta_{\boldsymbol{B}}=\frac{1}{\sqrt{1-2\cdot r(P)}}.$ (3.7)
@@
$\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$
@@
$\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}>\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$
(iii) Recall that $P_{m}$ is the regular $m$-gon. Theorem 1.1 (iii) shows that $r(P_{m})=r(P)$ if and only if $m$ is odd, and $P_{m}$ is uniquely optimal if and only if $m$ is a prime. According to Theorem 1.2, the matrix constructed from $P_{m}$ corresponds to the harmonic frame $\boldsymbol{E}_{m}$; in particular, $\boldsymbol{E}_{m}$ consists of the vectors $f^{-1}(\boldsymbol{e}_{j})$ obtained from the edge vectors $\boldsymbol{e}_{j}$ of $P_{m},j=1,\ldots,m$. Combining Theorem 1.2 and Theorem 1.1, we conclude that $\boldsymbol{E}_{m}$ has the minimal condition number if and only if $m$ is odd. Moreover, $\boldsymbol{E}_{m}$ is the unique matrix with minimal condition number if and only if $m$ is a prime. ∎